OpenStudy (anonymous):

Is the correct answer the last answer? 3 What is the value of the x variable in the solution to the following system of equations? 4x − 3y = 3 5x − 4y = 3 Choices: x can be any number as there are infinitely many solutions to this system There is no x value as there is no solution to this system −3 3

OpenStudy (phi):

if we multiply the top equation by 4 we get 16x -12y = 12 if we mult the bottom equation by -3 we get -15x +12y = -9 add these equations together x + 0 = 3 or x=3

OpenStudy (anonymous):

Ok that is what I got but wanted to make sure i did it correct. Thank you so much.